OpenStudy (anonymous):

x^2+4x-12=0 solve using completing the square

OpenStudy (zzr0ck3r):

x^2+4x-12=0 (x+2)^2 = 12 + 4 x+2 = +-sqrt(16) x = +-4-2 so x = -6 or x = -2

OpenStudy (anonymous):

-6 and -2

OpenStudy (anonymous):

-6 and 2

OpenStudy (anonymous):

x=4-2= 2 @zzr0ck3r

OpenStudy (zzr0ck3r):

yes 2 not -2 sorry